OpenStudy (anonymous):

MEDAL & FAN ANYBODY ? The volume of a sphere is 3,000 m.3. What is the surface area of the sphere to the nearest square meter? a. 2,158 m2 b. 37,699 m2 c. 165 m2 d. 1,018 m2

OpenStudy (anonymous):

Do you know the formula to calculate the surface area of the sphere and its volume?

OpenStudy (anonymous):

@tinybookworm no i don not .

OpenStudy (anonymous):

You can look up. But here, the surface area of sphere is \(S=4\pi r^2\) and its volume is \(\frac{4}{3} \pi r^3\).

OpenStudy (anonymous):

@tinybookworm well what should i do step by step ?

OpenStudy (anonymous):

Now can you calculate the radius based on the volume?

OpenStudy (anonymous):

im not sure how to do that @tinybookworm

OpenStudy (anonymous):

\(V= 3000= \frac{4}{3}\pi r^3\). Therefore, \(r^3=3000\times \frac{3}{4} \div \pi\), right?

OpenStudy (anonymous):

Which equals 6975 correct ? @tinybookworm

OpenStudy (anonymous):

Which one? r???

OpenStudy (anonymous):

im sorry , i got 716.56 @tinybookworm

OpenStudy (anonymous):

Correct. So \(r^3=716.56\), then r = ...

OpenStudy (anonymous):

would i divide that by 3 ? @tinybookworm

OpenStudy (anonymous):

No. \(r=\sqrt[3]{716.56}\)

OpenStudy (anonymous):

i dont know how to do that @tinybookworm

OpenStudy (anonymous):

Can your calculator do this one?

OpenStudy (anonymous):

i dont exactly have the technology @tinybookworm

OpenStudy (anonymous):

I can do it for you. So \(r=\sqrt[3]{716.56}\approx 8.95\). We have r, then we can calculate the surface area using the formular \(S=4\pi r^2\).

OpenStudy (anonymous):

@tinybookworm is it d ?

OpenStudy (anonymous):

i got it . thanks

OpenStudy (anonymous):

Yes. Good job. Hope you can solve this from now. :)